• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 11:48:17 on October23,2002

    Entry number 88870


    keyword=Controlled Access summary

    Ed, Keith, Mike, and Dave went into the hall during the MD down.

    Ed took a look at the magnets and saw no obvious problems.

    According to the target group, the fans in both loops were fixed "by magic" during the controlled access.

    It appears that they were able to take both fans up to 60 Hz while the heater power dropped by less than 100 W.

    Loop 2 fan is now operating at 60 Hz (25%), with the heater producing 570 W (no
    beam). The Loop 1 fan is back to 24 Hz (10%), with 125W from the heater.
    To my knowledge, the JT valves are unchanged from before.